OpenStudy (anonymous):

An urn contains 10 white balls and 5 blue balls. Draws are made repeatedly from the urn as follows. On each draw, a ball is drawn and its color noted; then it is replaced in the urn along with 3 more balls of its color. help . Find the chance that the second ball drawn is blue = 5/15*8/18 . is correct ?

OpenStudy (anonymous):

Find the chance that the first ball drawn is blue =1/3

OpenStudy (anonymous):

It's just \[\frac{ 8 }{ 18}\] for the chance of blue being drawn the 2nd time when 3 more blue balls were added. You are not finding the chance that a blue ball is drawn two times in a row, just the 2nd time.

OpenStudy (anonymous):

or 7/17 ?

OpenStudy (anonymous):

It's 18 I think, 15+3 = 18

OpenStudy (anonymous):

the denominator

OpenStudy (anonymous):

mm ok ..thk

OpenStudy (anonymous):

now is question Find the chance that the first ball drawn is blue, given that the second ball drawn is blue ? P(B1/B2)

OpenStudy (anonymous):

my aswer is P(B1/B2)=4/9 is correct ? .. look my diagram

OpenStudy (anonymous):

P(B1 and B2) =? hello ?

OpenStudy (anonymous):

is 4/27 ?

OpenStudy (anonymous):

P(B1/B2) is 4/9